# Understanding Oversmoothing in Diffusion-Based GNNs From the Perspective of Operator Semigroup Theory

### v1

@@ 6.3 Node Classification
Datasets. We evaluate the performance of our proposed method in comparison with existing GNN architectures. The focus is on both heterophilic and homophilic graph datasets to showcase the model’s versatility and robustness across diverse real-world scenarios. For heterophilic datasets, we utilize six datasets known for their low homophily ratios as identified in (Pei et al., 2020), including Chameleon and Squirrel (Rozemberczki et al., 2021), Film (Tang et al., 2009), Texas, Wisconsin, and Cornell datasets from the WebKB collection. For homophilic datasets, we employ Cora (McCallum et al., 2000), CiteSeer (Sen et al., 2008) and PubMed (Yang et al., 2016). For data splits, we adopt the methodology from (Pei et al., 2020), ensuring consistency and comparability in our evaluations. The performance is gauged in terms of accuracy, with both mean and standard deviation reported. Each experiment is conducted over 10 fixed train/validation/test splits to ensure the reliability and reproducibility of results.
@@
Baselines. Our model is benchmarked against a comprehensive set of GNN architectures, encompassing both traditional models: GCN (Kipf & Welling, 2017), GAT (Velickovic et al., 2017) and GraphSage (Hamilton et al., 2017), and recent ODE-based models: Continuous Graph Neural Networks (CGNN) (Xhonneux et al., 2020), GRAND (Chamberlain et al., 2021b), Sheaf (Bodnar et al., 2022), GRAFF (Di Giovanni et al., 2022) and GREAD (Choi et al., 2023).
@@
Results. The data presented in Table 1 illuminates the comparative efficacy of diverse graph neural network models when applied to datasets characterized by varying degrees of homophily. Notably, the diffusion-based models, augmented with our proposed ergodicity breaking terms, demonstrate remarkable robustness and adaptability, emerging as the top-performing models on 5 out of 9 datasets. Furthermore, they secure the second position on 7 out of 9 datasets, consistently achieving top-three performance overall. Particularly noteworthy is their exceptional performance on datasets with a high homophily level, such as Citeseer, Pubmed, and Cora, as well as on the Texas dataset, which exhibits the lowest level of homophily. The ‘Ours’ suite of models demonstrates commendable performance by attaining the highest ranks across all these datasets. The ‘Ours’ suite of models demonstrates commendable performance by attaining the highest ranks across all these datasets.
7 Conclusions
@@
Broader Impact
Our contributions to the field of diffusion-based Graph Neural Networks (GNNs) have broad implications for both theoretical research and practical applications in data science and artificial intelligence. By introducing a novel framework grounded in operator semigroup theory, we address the pervasive issue of oversmoothing, thereby enhancing the effectiveness and interpretability of GNNs. The introduction of an ergodicity-breaking condition, validated by both synthetic and real-world data, extends the utility of GNNs beyond traditional boundaries, offering new avenues for analysis and application. Our probabilistic interpretation further bridges theoretical gaps, fostering a deeper understanding of GNN dynamics. This work not only advances the state-of-the-art in GNN methodologies but also expands their applicability in solving complex networked problems, marking a significant step forward in the development of more robust and versatile AI systems.
@@
Lemma A.1 (Theorem 2.1.2 of (Norris, 1998) ).
@@
Proof of proposition 3.2.
Lemma A.1 demonstrated that $P_{t}$ is a stochastic matrix adhere to: $0\leq p_{ij}<\infty$ , $\sum_{j\in I}p_{ij}=1$ for all $i$, leading to (i) and (ii). ∎
A.2 Proof of Theorem 4.3
Lemma A.2.
@@
Suppose $\mu$ is an eigenvector of $A^{T}$ with eigenvalue 1, then $A^{T}\mu=\mu$ and $\mu(v)=\sum_{u}a(u,v)\mu(u)$
@@
We first consider operator $\mathcal{A}$ define in Eq.(6) is self-adjoint, that is, $\mathcal{A}$ is equal to its adjoint operator $\mathcal{A}^{*}$. To be specific, $\langle\mathcal{A}f,g\rangle=\langle f,\mathcal{A}g\rangle$, for every $f,g\in C(\mathcal{V})$, where inner product $\langle f,g\rangle:=\int_{\mathcal{V}}fg\mathrm{d}\mu$.
Proof of Theorem 4.3.
Since $\mathcal{G}$ is a connected graph, then $A=(a(u,v))_{u,v\in\mathcal{V}}$ is irreducible, that is, for every pair of vertices $(u,v)\in\mathcal{V}\times\mathcal{V}$, there exists a path $(u=u_{0},u_{1},\ldots,u_{k-1},u_{k}=v)$ in $\mathcal{V}$, such that for $i=0,1,\ldots,k-1$, $a(u_{i},u_{i+1})>0$. Equivalently, $A$ is not similar via a permutation to a block upper triangular matrix. From Perron-Frobenius theorem for irreducible matrices, we know that the maximal eigenvalue $\lambda_{0}=0$ of $\mathcal{A}$ is unique (algebra multiplicity is $1$) and the geometric multiplicity of $\lambda_{0}$ is $1$, that is, the corresponding normalized eigenfunction of $\lambda_{0}$ is unique, noted as $f$. Consider
@@
$=\frac{1}{2}\left[\sum_{u\in\mathcal{V}}\sum_{v\in\mathcal{V}}a(u,v)(f(v)-f(u))f(u)\mu(u)+\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{V}}a(v,u)(f(u)-f(v))f(v)\mu(v)\right]$
$=\frac{1}{2}\sum_{u\in\mathcal{V}}\sum_{v\in\mathcal{V}}a(u,v)\left[(f(v)-f(u))f(u)+(f(u)-f(v))f(v)\right]\mu(u)$
@@
$f$ is the corresponding normalized eigenfunction of $\lambda_{0}=0$, that is, $\mathcal{A}f=0$, then $\langle\mathcal{A}f,f\rangle=0$. Therefore
$f(u)=f(v):=c,\quad\forall u,v\in\mathcal{V}.$
Notice the uniqueness of $f$, operator $\mathcal{A}$ is ergodic. ∎
A.3 Proof of Theorem 4.4
@@
where $\lambda_{k},k=0,1,\ldots,N$ are the eigenvalues of $\mathcal{A}$, $e_{\lambda_{k}}$ are corresponding normalized eigenfunctions of $\lambda_{k}$, $E_{\lambda_{k}}$ are corresponding projection operators of $\lambda_{k}$, and $f_{\lambda_{k}}$ is the projection of $f$ related to $e_{\lambda_{k}}$ For $t\geq 0$,
$P_{t}f=e^{t\mathcal{A}}f=\sum_{k=0}^{N}e^{\lambda_{k}t}f_{\lambda_{k}}e_{\lambda_{k}}.$
Since graph $\mathcal{G}$ is connected, Perron-Frobenius theorem implies that eigenvalues of $\mathcal{A}$ satisfy $0=\lambda_{0}>\lambda_{1},\ldots,\lambda_{N}$. As $t\rightarrow\infty$,
$P_{t}f(u)\rightarrow f_{\lambda_{0}}e_{\lambda_{0}}.$
Since operator $\mathcal{A}$ is ergodic, $f_{\lambda_{0}}e_{\lambda_{0}}$ is a constant. Notice
@@
If $A$ is not symmetric, while for every pair of vertices $(u,v)\in\mathcal{V}\times\mathcal{V}$,
$\mu(u)a(u,v)=\mu(v)a(v,u),$
which is so called detailed balance condition, we say $\mu$ is reversible. We can simply symmetrize $A$ by
$\tilde{a}(u,v):=\sqrt{\mu(u)}a(u,v)\frac{1}{\sqrt{\mu(v)}}.$
Consider the symmertrized operator $\tilde{\mathcal{A}}$ satisfying : $\tilde{\mathcal{A}}f(u)=\sum_{v\in\mathcal{V}}\tilde{a}(u,v)(f(v)-f(u))$, we can obtain similar conclusions.
A.4 Proof of Theorem 4.5
@@
Since $\mathcal{A}b=\sum_{v\in\mathcal{V}}a(u,v)(b-b)=0$, $\mathcal{C}b\neq 0$. It is contradict to $(\mathcal{A}+\mathcal{C})f_{c}(u)=0$. Therefore $\mathcal{A}+\mathcal{C}$ is not ergodic. ∎
@@
$\mathcal{A}h_{i}(u,t)$ $=\lim\limits_{r\rightarrow 0^{+}}\frac{P_{r}h_{i}(u,t)-h_{i}(u,t)}{r}$
$=\lim\limits_{r\rightarrow 0^{+}}\frac{\mathbb{E}^{u}[h_{i}(\mathbf{X}_{r},t)]-h_{i}(u,t)}{r}$
@@
Let $\{\mathscr{F}_{t}\}$ be the right continuous $\sigma$-algebra filtration with respect to which $\mathbf{X}_{t}$ is adapted. From Chapman–Kolmogorov equation (or Markov property),
@@
$=\lim\limits_{r\rightarrow 0^{+}}\frac{h_{i}(u,t+r)-h_{i}(u,t)}{r}$
$=\frac{\partial h_{i}(u,t)}{\partial t}.$

### v2

@@ 7.3. Node Classification
Datasets. We evaluate the performance of our proposed method in comparison with existing GNN architectures. The focus is on both heterophilic and homophilic graph datasets to showcase the model’s versatility and robustness across diverse real-world scenarios. For heterophilic datasets, we utilize six datasets known for their low homophily ratios as identified in (Pei et al., 2020), including Chameleon and Squirrel (Rozemberczki et al., 2021), Film (Tang et al., 2009), Texas, Wisconsin, and Cornell datasets from the WebKB collection. For homophilic datasets, we employ Cora (McCallum et al., 2000), CiteSeer (Sen et al., 2008) and PubMed (Yang et al., 2016). For data splits, we adopt the methodology from (Pei et al., 2020), ensuring consistency and comparability in our evaluations. The performance is gauged in terms of accuracy, with both mean and standard deviation reported. Each experiment is conducted over 10 fixed train/validation/test splits to ensure the reliability and reproducibility of results. In addition, following the settings in GRAND (Chamberlain et al., 2021b), additional datasets include the coauthor graph CoauthorCS (McAuley et al., 2015), the Amazon co-purchasing graphs Computer and Photo (McAuley et al., 2015), and the OGB Arxiv dataset (Hu et al., 2020).
@@
Baselines. Our model is benchmarked against a comprehensive set of GNN architectures, encompassing both traditional models: PairNorm (Zhao and Akoglu, 2020), GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018) and GraphSage (Hamilton et al., 2017), and recent ODE-based models: Continuous Graph Neural Networks (CGNN) (Xhonneux et al., 2020), GRAND (Chamberlain et al., 2021b), GRAFF (Di Giovanni et al., 2022) and GREAD (Choi et al., 2023).
Hardwares. All experiments reported in this work were conducted within a uniform computational environment to ensure reproducibility and consistency of results. The specifications of the environment are as follows: Operating System: Ubuntu 18.04 LTS; Programming Language: Python 3.10.4; Deep Learning Framework: PyTorch 2.0.1; Graph Neural Network Library: PyTorch Geometric 2.4.0; Differential Equation Solver: TorchDiffEq 0.2.3; GPU Computing: CUDA 11.7; Processor: AMD EPYC 7542 32-Core Processor; Graphics Card: NVIDIA RTX 3090.
@@
Results. The data presented in Table 1 illuminates the comparative efficacy of diverse graph neural network models when applied to datasets characterized by varying degrees of homophily. Notably, the diffusion-based models, augmented with our proposed ergodicity breaking terms, demonstrate remarkable robustness and adaptability, emerging as the top-performing models on 8 out of 9 datasets. Furthermore, they consistently secure the top-three performance overall. Particularly noteworthy is their exceptional performance on datasets with low homophily levels, such as Wisconsin and Cornell, as well as on the Pubmed dataset, which exhibits a high level of homophily. For the additional results presented in Table 2, ’Ours-Expn’ achieves the highest accuracy on the Computers, Photo, and CoauthorCS datasets, while ’Ours-Log’ achieves the highest accuracy on the ogbn-arxiv dataset. These findings indicate that the proposed breaking term is effective in enhancing performance for node classification tasks.
7.4. Discussion on Killing Process
In this part, we provide a detailed discussion of the killing process introduced in Section 5.3. According to Theorem 3, the killing process follows the ODE dynamics described as: $\frac{\partial H(t)}{\partial t}=[A(H(t))-I]H(t)-c(H(t))H(t)$, from which the following observations can be made: (1) Unlike the breaking term introduced in Section 4.3, here $c(H(t))$ is negative (element-wise in the vector sense) and can depend solely on the node features $H(t)$, independent of the graph structure; (2) From the proof provided in Appendix B.3, the node feature under the killing process can be expressed as $h(X_{t})=\mathbb{E}^{u}[f(X_{t})e^{-\int_{0}^{t}c(X_{s})\mathbf{d}s}]$. This highlights that the parameter $c(H(t))$ in the killing process directly controls the convergence rate of the node features. To more clearly illustrate the impact of the design of $c(H(t))$ on oversmoothing, we consider a special case where c is a constant. In this scenario, the node features are given by $h(X_{t})=\mathbb{E}^{u}[f(X_{t})e^{-ct}]$. Under this setting, nodes undergo oversmoothing at an exponential rate, with the rate controlled by the magnitude of $c$. We employ the log-scale Dirichlet energy to represent the extent of oversmoothing in diffusion-based GNNs as the number of layers increases. The results are presented in Table 3.
Table 3. Dirichlet energy in log scale as layers increase under different $c$ values.
Layers
c 1 10 20 60 80 90 100
1e-3 4.66 -0.80 -2.15 -4.39 -5.09 -5.39 -5.66
1e-2 4.66 -0.96 -2.49 -5.47 -6.52 -7.00 -7.46
1e-1 4.66 -2.60 -5.95 -16.19 -20.73 -21.06 -24.78
5e-1 4.66 -9.87 -21.23 -21.33 -21.96 -26.70 -23.96
1 4.66 -18.96 -25.00 -22.24 -21.43 -23.00 -21.63
It can be observed that as $c$ increases, the rate of oversmoothing accelerates significantly. In practice, the gating mechanism introduced in (Rusch et al., 2023) can be regarded as a learning-based approach to determine $c(H(t))$, thereby aiming to alleviate the problem of oversmoothing. We speculate that designing different forms of $c(H(t))$ can lead to varying effects both in theory and practice. Our work aims solely to establish this connection, while the more theoretical design of $c(H(t))$ is left as future work.
7.5. Effects of Truncated Orders $N$
For two datasets with distinct homophily ratios, Texas and Cora, we present in Table 4 the impact of different Truncated Orders on the performance of the node classification task. Results reveal that the performance is significantly impacted by the interplay between homophily and $N$. In low-homophily settings (Texas), performance is more sensitive to $N$, with Ours-Expn exhibiting superior robustness, likely due to its ability to filter out less relevant information from multi-hop aggregation. Conversely, in high-homophily settings (Cora), performance is less sensitive to changes in $N$, as immediate neighborhood information is often sufficient for classification, with all models exhibiting similar performance across values of $N$. These findings highlight the necessity of tailoring the truncation order $N$ to the dataset’s homophily level to optimize GNN performance.
Table 4. Node Classification Results w.r.t. Truncated Orders
Texas Cora
Truncated Orders N=1 N=2 N=3 N=1 N=2 N=3
Ours-Exp 84.06 79.73 80.27 87.04 87.97 87.30
Ours-Expn 88.65 85.41 87.03 88.31 88.07 87.85
Ours-Log 86.49 87.03 85.95 87.40 88.07 87.44
8. Conclusions and Limitations
@@
While we discuss the nonlinear graph diffusion operator and show that attention-based nonlinear graph diffusion still suffers from oversmoothing, quantitative analysis in the continuous case is still difficult. In our future work, we aim to investigate the specific designs of the breaking term and the killing process to better understand their precise effects on node-level features, and will explore oversmoothing in the nonlinear case using advanced mathematical tools such as nonlinear operator semigroups. Moreover, our experimental results primarily focus on node classification tasks, which are well-suited for evaluating whether a model is influenced by oversmoothing. Exploring the impact of the oversmoothing problem on other types of tasks, such as graph-level tasks, could serve as a promising direction for future research.
Acknowledgments
This work is supported by National Key R&D Program of China (No.2021YFA1000403), National Natural Science Foundation of China (No.12401666, 11991022, U23B2012, 12326611) and the Fundamental Research Funds for the Central Universities, Nankai University (No.054-63241437).
@@ A.1. Proof of Proposition 1
Lemma 1 (Theorem 2.1.2 of (Norris, 1998) ).
@@
Lemma 1 demonstrated that $P_{t}$ is a stochastic matrix adhere to: $0\leq p_{ij}<\infty$ , $\sum_{j\in I}p_{ij}=1$ for all $i$, leading to (i) and (ii).
A.2. Proof of Theorem 3
Lemma 2.
@@ A.2. Proof of Theorem 3
Suppose $\mu$ is an eigenvector of $A^{T}$ with eigenvalue 1, then $A^{T}\mu=\mu$ and $\mu(v)=\sum_{u}a(u,v)\mu(u)$,
@@
We first consider operator $\mathcal{A}$ define in Eq.(7) is self-adjoint. To be specific, $\langle\mathcal{A}f,g\rangle=\langle f,\mathcal{A}g\rangle$, for every $f,g\in C(\mathcal{V})$, where inner product $\langle f,g\rangle:=\int_{\mathcal{V}}fg\mathrm{d}\mu$.
Proof of Theorem 3.
Since $\mathcal{G}$ is a connected graph, then $A=(a(u,v))_{u,v\in\mathcal{V}}$ is irreducible, that is, for every pair of vertices $(u,v)\in\mathcal{V}\times\mathcal{V}$, there exists a path $(u=u_{0},u_{1},\ldots,u_{k-1},u_{k}=v)$ in $\mathcal{V}$, such that for $i=0,1,\ldots,k-1$, $a(u_{i},u_{i+1})>0$. Equivalently, $A$ is not similar via a permutation to a block upper triangular matrix.
From Perron-Frobenius theorem for irreducible matrices, we know that the maximal eigenvalue $\lambda_{0}=0$ of $\mathcal{A}$ is unique (algebra multiplicity is $1$) and the geometric multiplicity of $\lambda_{0}$ is $1$, that is, the corresponding normalized eigenfunction of $\lambda_{0}$ is unique, noted as $f$. Consider
@@
$=\frac{1}{2}[\sum_{u\in\mathcal{V}}$ $\sum_{v\in\mathcal{V}}a(u,v)(f(v)-f(u))f(u)\mu(u)$
$+\sum_{v\in\mathcal{V}}\sum_{u\in\mathcal{V}}a(v,u)(f(u)-f(v))f(v)\mu(v)]$
@@
$f$ is the corresponding normalized eigenfunction of $\lambda_{0}=0$, that is, $\mathcal{A}f=0$, then $\langle\mathcal{A}f,f\rangle=0$. Therefore $f(u)=f(v):=c,\;\forall u,v\in\mathcal{V}.$ Notice the uniqueness of $f$, operator $\mathcal{A}$ is ergodic. ∎
A.3. Proof of Theorem 4
@@
where $\lambda_{k},k=0,1,\ldots,N$ are the eigenvalues of $\mathcal{A}$, $e_{\lambda_{k}}$ are corresponding normalized eigenfunctions of $\lambda_{k}$, $E_{\lambda_{k}}$ are corresponding projection operators of $\lambda_{k}$, and $f_{\lambda_{k}}$ is the projection of $f$ related to $e_{\lambda_{k}}$ For $t\geq 0$, $P_{t}f=e^{t\mathcal{A}}f=\sum_{k=0}^{N}e^{\lambda_{k}t}f_{\lambda_{k}}e_{\lambda_{k}}.$ Since graph $\mathcal{G}$ is connected, Perron-Frobenius theorem implies that eigenvalues of $\mathcal{A}$ satisfy $0=\lambda_{0}>\lambda_{1},\ldots,\lambda_{N}$. As $t\rightarrow\infty$, $P_{t}f(u)\rightarrow f_{\lambda_{0}}e_{\lambda_{0}}.$ Since operator $\mathcal{A}$ is ergodic, $f_{\lambda_{0}}e_{\lambda_{0}}$ is a constant. Notice
@@
If $A$ is not symmetric, while for every pair of vertices $(u,v)\in\mathcal{V}\times\mathcal{V}$, $\mu(u)a(u,v)=\mu(v)a(v,u),$ which is so called detailed balance condition, we say $\mu$ is reversible. We can simply symmetrize $A$ by $\tilde{a}(u,v):=\sqrt{\mu(u)}a(u,v)\frac{1}{\sqrt{\mu(v)}}.$ Consider the symmertrized operator $\tilde{\mathcal{A}}$ satisfying : $\tilde{\mathcal{A}}f(u)=\sum_{v\in\mathcal{V}}\tilde{a}(u,v)(f(v)-f(u)),$ we can obtain similar conclusions.
A.4. Proofs for Theorem 6
Proof.
For simplicity, we denote the attention matrix $A(H(t))$ defined in (3) as $A^{(t)}$, satisfies
$\begin{cases}A^{(t)}(u,v)=\frac{e^{b^{(t)}_{u,v}}}{\sum_{w\in\mathcal{N}(u)}e^{b^{(t)}_{u,w}}}&\text{if}(u,v)\in\mathcal{E},\\ A^{(t)}(u,v)=0&\text{if}(u,v)\notin\mathcal{E}.\end{cases}$
where $b^{(t)}_{u,v}=\Phi(h_{t}(u),h_{t}(v)),\;(u,v)\in\mathcal{E}$. Through the definition of $H_{t+1}$, we have $\|H_{t+1}\|_{max}\leq\|A^{(t)}A^{(t-1)}...A^{(0)}\|_{max}\|H_{0}\|_{max}$, where $\|\cdot\|_{max}$ denotes the matrix max norm, i.e., let $E=(e_{ij})\in\mathbb{R}^{m\times n}$, $\|E\|_{max}=\max\limits_{i,j}e_{ij}$. Since $A^{(t)}(u,v)\in[0,1]$ for all $t$ and all $(u,v)$, $\|H_{t+1}\|_{max}\leq\|H_{0}\|_{max}$ Hence there exist $C_{0}\geq 0$ such that for $t\in\mathbb{N}$, $\|H_{t+1}\|_{max}\leq C_{0}.$ The continuous assumption of attention function $\Phi$ implies there exist $C\geq 0$ such that $0\leq b^{(t)}_{u,v}\leq C$ for all $t$, we thus obtain that
$A^{(t)}(u,v)=\frac{e^{b^{(t)}_{u,v}}}{\sum_{w\in\mathcal{N}(u)}e^{b^{(t)}_{u,w}}}\geq\frac{1}{Ne^{c}},$
for all $(u,v)\in\mathcal{E}$. Hence there exist $\epsilon>0$ such that $A^{(t)}(u,v)\geq\epsilon$ for all $t\in\mathbb{N}$ and $(u,v)\in\mathcal{E}$, impling that $A^{(t)}(u,v)$ satisfies the Doeblin’s condition (Saloff-Coste and Zúñiga, 2009), yielding
$max\{\|P_{t}(u,\cdot)-P_{t}(v,\cdot)\|_{TV}\}\leq(1-\epsilon)^{t}\rightarrow 0,\quad t\rightarrow\infty$
completing the proof. ∎
From Theorem 6 we can derive another quantitative lower bounds for $P_{t}$ using mathematical Induction. To be specific, we aim at proving the following statement: for all $t\geq 0$ and all $(u,v)\in\mathcal{E}$, $P_{t}(u,v)\geq\epsilon^{t+1}$. When $t=0$, $P_{0}=A^{(0)}$, it is obvious to see the statement holds. We assume the statement holds at step $t-1$, implying that $P_{t-1}(u,v)\geq\epsilon^{t}$. The self-loop assumption indicate that
$P_{t}(u,v)\geq A^{(t)}(u,u)P_{t-1}(u,v)$
[…]
A.5. Proof of Theorem 7
@@
Since $\mathcal{A}b=\sum_{v\in\mathcal{V}}a(u,v)(b-b)=0$, $\mathcal{C}b\neq 0$. It is contradict to $(\mathcal{A}+\mathcal{C})f_{c}(u)=0$. Therefore $\mathcal{A}+\mathcal{C}$ is not ergodic.
@@ B.1. Proof of Theorem 1
$\mathcal{A}h_{i}(u,t)$ $=\lim\limits_{r\rightarrow 0^{+}}\frac{\mathbb{E}^{u}[h_{i}(\mathbf{X}_{r},t)]-h_{i}(u,t)}{r}$
@@
Let $\{\mathscr{F}_{t}\}$ be the right continuous $\sigma$-algebra filtration with respect to which $\mathbf{X}_{t}$ is adapted. From Chapman–Kolmogorov equation,
@@
$=\lim\limits_{r\rightarrow 0^{+}}\frac{h_{i}(u,t+r)-h_{i}(u,t)}{r}=\frac{\partial h_{i}(u,t)}{\partial t}.$